OpenStudy (anonymous):

Under water, pressure increases 4.3 pounds per square inch (psi) for every 10 feet you descend. This is represented by the equation p = 0.43d + 14.7, where p is the psi and d is the depth in ft. Find the pressure at 200 feet. (Points : 2) 215.13 psi 860 psi 100.7 psi

OpenStudy (danjs):

p(200) =

OpenStudy (anonymous):

p = 0.43d + 14.7 p = 0.43 X (200) + 14.7 p = 86 + 14.7 p = 100.7

OpenStudy (anonymous):

@DanJS Right?

OpenStudy (danjs):

yep

OpenStudy (anonymous):

Perfect thanks:)

OpenStudy (danjs):

Don't know, they might have a question, What is the pressure from the Air or something, P(0)= 14.7 psi , at sea level

OpenStudy (anonymous):

oh okay